# Latency Adjustable Transformer Encoder for Language Understanding

### v8

@@ Abstract
Adjusting the latency, power, and accuracy of natural language understanding models is a desirable objective of an efficient architecture. This paper proposes an efficient Transformer architecture that adjusts the inference computational cost adaptively with a desired inference latency speedup. In fine-tuning phase, the proposed method detects less important hidden sequence elements (word-vectors) and eliminates them in each encoder layer using a proposed Attention Context Contribution (ACC) metric. After the fine-tuning phase, with the novel offline-tuning property, the inference latency of the model can be adjusted in a wide range of inference speedup selections without any further training. The proposed method is applied to the BERTbase, GPT-2 and Flan-T5 models for evaluation. Extensive experiments show that most of the word-vectors in higher Transformer layers have less contribution to the subsequent layers; hence, they can be eliminated to improve the inference latency. Experimental results on extensive sentiment analysis, classification, text generation tasks and regression benchmarks like GLUE showed that the method is effective in various datasets with minimal impact on the input’s global context. The method was also evaluated under the instruction tuning paradigm, and its performance was measured using different types of prompting. The proposed method mathematically and experimentally improves the inference latency of BERTbase and GPT-2 by up to 4.8 and 3.72 times with less than 0.75% accuracy drop and passable perplexity on average. The suggested approach posits that in Large Language Models (LLMs), although the complete network is necessary for training, it can be truncated during the fine-tuning phase.
@@ I Introduction
Recently, Transformer [1] based architectures achieved remarkable success in various Natural Language Processing (NLP) tasks. However, the training difficulty, inference latency, and small size of available datasets are the main concerns with using these models in real applications [2]. Fine-tuning a pre-trained language model on downstream tasks mitigates some of these concerns, including training effort and dataset size. GPT [3], BERT [4], and XLNet [5] are examples of recent state-of-the-art pre-trained language models. The number of parameters in these models is more than hundreds of millions. Hence, the computational complexity bears billions of Floating Point Operations (FLOPs) for processing a single sample. In some cases, the inference phase computational effort is beyond the ability of resource-limited systems such as edge devices. Therefore, baseline Transformer models cause unacceptable latency and energy usage limitations in such devices.
One of the most effective pre-trained language models is Bidirectional Encoder Representations from Transformers (BERT) [4], which has multi-layer Transformer-based architecture. BERTbase consists of an embedding layer and 12 stacked encoder layers. Each encoder contains an attention layer with 12 heads and multiple fully connected layers. There are about 110M parameters in BERTbase model, the pre-training of which has taken 79 hours on 64 Tesla V100 GPUs [2].
@@
This study introduces an innovative attention layer designed to enhance the efficiency of inference stages in Transformer-based models with an adjustable approach. The suggested method involves a sorting and elimination process for hidden sequence elements (word-vectors), resulting in a reduction of effective word-vectors in each layer and a decrease in the number of Floating Point Operations (FLOPs). The primary contributions of this research can be summarized as follows:
@@
  • An accurate analytical inference speedup estimation is presented for the proposed method that helps a system designer determine the hyper-parameter selection at the fine-tuning phase.
@@
  • The proposed approach can be applied to diverse encoder and decoder Transformer-based architectures and preserves the input’s global context.
@@ II-B Head/Channel Pruning
A parameter reduction method was proposed by identifying the most critical heads in each encoder layer using layer-wise relevance propagation and pruning the redundant heads [32]. It showed that the earlier encoder layer’s heads are much more critical than the last. Moreover, in [33], it is shown that around 17% of attention heads can be removed at the test time without significantly impacting inference performance. In [34], a rapid post-training pruning framework tailored for Transformers is introduced. This framework involves three key fine-tuning stages: mask search, mask rearrangement, and mask tuning. Notably, the method obviates the necessity for any retraining, presenting an efficient approach to achieve model compression in the Transformer architecture.
@@ II-E Non-Pruning Methods
ALBERT is a lighter version of BERT model [45]. It uses cross-layer weight sharing and a modified embedding layer structure, making the embedding size independent of the encoder’s hidden size. Moreover, the new Sentence Order Prediction (SOP) task is utilized in the pre-training phase instead of the Next Sentence Prediction task. ALBERT reports state-of-the-art results with its parameter reduction and modified pre-training policy. ALBERT is nine times smaller and 1.2 times faster than BERTbase, with less than 2.5% on average accuracy drop. ALBERT is an example of the parameter-sharing technique.
DistilBERT [46], TinyBERT [47] and MobileBERT [48] use the knowledge distillation technique during the pre-training and offer up to 5.5 times faster inference on downstream tasks than BERTbase. In [49] a Self-Distillation (SD) mechanism obtains high-accuracy models directly without going through an assistive model. The main concerns with these methods are the pre-training effort of the model from scratch, extra training steps, and significant accuracy degradation.
In [50], the authors propose sufficient conditions under which they prove that a sparse attention model can universally approximate any sequence-to-sequence function. They show that sparse Transformers with only $\mathcal{O}_{n}$ connections per attention layer can approximate the same function class as the dense model with $n^{2}$ connections. Consequently, in recent years, many works [51, 52, 53, 54, 55, 56, 57] have proposed methods that aim to reduce the complexity of the attention mechanism from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$, where $N$ is the sequence length.
@@
The previous works are trying to reduce the model size and inference latency. These works show that significant part computations in the models appear unnecessary, and models with less computational complexity can be obtained from the original ones. The main drawbacks of these methods are pre-training from scratch, multi-phase fine-tuning, and significant inference accuracy drop. In these methods, the inference speedup is set in the fine-tuning phase, and the achieved speedup is fixed in the inference phase. The proposed method tries to mitigate the previous drawbacks. It presents the powerful offline-tuning mechanism that controls the inference speedup of the model after the fine-tuning phase.
The impact of the speedup methods on the model stability and the input’s global context understanding is an important concern that must be studied in methods evaluations. Backpack learns multiple non-contextual sense vectors for each word in a vocabulary and represents a word in a sequence as a context-dependent, non-negative linear combination of sense vectors in this sequence [61]. The Backpack Model shows that each word in the input sequence can interpreted as a sense vector that encodes different aspects of a word which is very sensitive. A good practice for assessing the performance of a method is evaluating it in a more complex task, like text generation tasks that inspect the model proficiency. Therefore the proposed method is applied to the GPT-2 model beside the BERTbase model and evaluated on several text generation tasks.
@@ III Proposed Method
This Section presents details of the proposed method and the analysis and reasoning behind the method. Section III-A presents the preliminary background of the BERTbase architecture, and Section III-B analyzes the FLOPs in the BERTbase model. The proposed encoder layer is based on an insight into the amount of the contribution of word-vectors at different layers, which is considered in Section III-C. In Section III-D, the proposed architecture is investigated in detail. In Section III-E, the proposed method is extended to decoder architecture and applied to the GPT-2 model.
@@ III-A Background
Fig. 1 shows BERTbase model architecture and output classifier. The input sentence goes through the tokenizer layer and embedding layer. Then the context is fed into 12 cascaded encoder layers to be processed. The pooler and classifier layers are added in the last layer to perform proper output for the desired task. In the figure, $T$ and $H$ are the numbers of input word-vectors and the hidden state vector size, respectively. Each encoder consists of a Bidirectional Multi-Head Self-Attention layer and a Feed-Forward layer with several tensor multiplications. The residual branches and normalization layers increase the training stability.
Figure 1: BERTbase [4] architecture with output layers
III-B BERT FLOPs Analysis
Table I shows the estimated FLOPs of BERTbase at different layers. The amount of FLOPs is estimated based on the number of input tokens ($T$), the number of encoder layers ($L$), the hidden state size ($H$), and the number of output labels ($N$).
TABLE I: BERTbase FLOPs Study
Layer Sub-layer FLOPs Estimationa FLOPs Estimation FLOPs Shareb Latency Sharec
Embedding Word-embeddings 0
Position-embeddings 0 $7TH$ 0.003% 0.2%
Layer-normalization $7TH$
Encoder (12 layers) Attention-self $6LTH^{2}$ + $2HTL^{2}$ $6LTH^{2}$ + $2HTL^{2}$ 25.1%
Attention-output $2LTH^{2}$ 99.8%
Intermediate $8LTH^{2}$ $18LTH^{2}$ 74.9%
Output $8LTH^{2}$
Classifier Pooler $2H^{2}$ $2H^{2}$ 0.0013% 0.036%
Output $2HN$
a $T$ = number of input tokens, $L$ = number of encoder layers, $H$ = hidden state size and $N$ = number of output labels
b With $T$ = 512, $L$ = 12, $H$ = 768 and $N$ = 2
c IMDB [62] dataset average relative inference time latency.
[…]
The inference latency and computational effort are directly related to the number of FLOPs. As a result, reducing the number of FLOPs at the encoder layers reduces the total number of FLOPs and inference latency.
@@ III-C Attention Context Contribution (ACC) metric
Fig. 2 illustrates the calculation of the Score Vector based on the attention probability matrix. Fig. 2 (left) shows the attention probability matrix of BERTbase model. The rows of this matrix are normalized to one. By averaging the attention probability matrix over the heads, the aggregated probability matrix is resulted, Fig. 2 (center). The SV results from the sum of this matrix along the unnormalized dimension as shown in Fig. 2 (right). The SV elements with higher score values correspond to the word-vectors that contribute more to the layer output.
@@
Fig. 3 illustrates BERTbase SV results on the IMDB sentiment analysis task[62] for an input sample. As shown in earlier encoder layers, the SV values approximately have uniform distributions, and in the last encoder layers, the distribution of SV values changes to approximately $\delta$ distribution. The distribution of SV values in the last layers indicates that the concentration of scores is high on a few word-vectors and others have less contribution to the self-attention layer output.
Figure 3: BERTbase[4] Score Vector (SV) results on the IMDB sentiment analysis task[62]
The SV elements are positive and with an average of one. So, the median of SV is bounded. Suppose the median of the SV is lower than its average. In that case, more than half of the elements are lower than the average value, and only a few are higher than the average value. It suggests that most of the word-vectors do not bear enough contribution to the layer output, and only a few word-vectors influence the final result of the layer. The median of the Score Vector is interpreted as the proposed Attention Context Contribution metric.
The ACC metric calculation is explained in Algorithm. 1. The algorithm inputs are BERT encoder word-vectors and the model parameters. The algorithm output is the ACC metric vector that suggests the contribution of the word-vectors in each encoder self-attention layer. For encoder layer $l$, the first step is to prepare the pre-computed self-attention probability matrix; $M_{Att}^{l}$. Then the average of this matrix over its heads is computed; $M_{H}^{l}$. In the next step, the resulting matrix’s summation over its unnormalized dimension is computed; $M_{S}^{l}$. This matrix is the Score Vector. Attention Context Contribution of the layer ($E_{ACC}^{l}$) is the median of the Score Vector.
@@
Fig. 4 shows the ACC metric for BERTbase encoder layers on IMDB [62] dataset and several General Language Understanding Evaluation (GLUE) [63] benchmark tasks. The fitted curve to the ACC metric results shows that the ACC metric is reduced at later layers gradually. It indicates that the fraction of the word-vectors that contribute more in the last layers is less in those layers; hence more word-vectors can be eliminated. By eliminating less important word-vectors in each layer, the processed word-vectors are decreased gradually, and the inference time of the model is decreased.
@@
Figure 4: BERTbase [4] Encoder Attention Context Contribution (ACC) metric results (blue bars) and second degree fitted curves (red lines) on SST-2 [64], MRPC [65], QQP [66], QNLI [67], RTE [68], MNLI [69], STS-B [70] and IMDB [62] tasks.

### v9

@@ Abstract
Adjusting the latency, power, and accuracy of natural language understanding models is a desirable objective of an efficient architecture. This paper proposes an efficient Transformer architecture that adjusts the inference computational cost adaptively with a desired inference latency speedup. In fine-tuning phase, the proposed method detects less important hidden sequence elements (word-vectors) and eliminates them in each encoder layer using a proposed Attention Context Contribution (ACC) metric. After the fine-tuning phase, with the novel offline-tuning property, the inference latency of the model can be adjusted in a wide range of inference speedup selections without any further training. Extensive experiments reveal that most word-vectors in higher Transformer layers contribute less to subsequent layers, allowing their removal to improve inference latency. Experimental results on various language understanding, text generation, and instruction tuning tasks and benchmarks demonstrate the approach’s effectiveness across diverse datasets, with minimal impact on the input’s global context. The technique improves Time-to-First-Token (TTFT) of Llama3 by up to 2.9x, with minor performance drop. The suggested approach posits that in Large Language Models (LLMs), although the complete network is necessary for training, it can be truncated during the fine-tuning phase.
@@ I Introduction
Recently, Transformer [1] based architectures achieved remarkable success in various Natural Language Processing (NLP) tasks. However, the training difficulty, inference latency, and small size of available datasets are the main concerns with using these models in real applications [2]. Fine-tuning a pre-trained language model on downstream tasks mitigates some of these concerns, including training effort and dataset size. BERT [3], XLNet [4], GPT [5], Llama [6], and Gemma [7] represent a spectrum of pre-trained language models, ranging from earlier developments to the latest advancements. These models vary significantly in size, ranging from hundreds of millions to hundreds of billions of parameters. Hence, the computational complexity bears trillions of Floating Point Operations (FLOPs) for processing a single sample. In some cases, the inference phase computational effort is beyond the ability of resource-limited systems such as edge devices. Therefore, baseline Transformer models cause unacceptable latency and energy usage limitations in such devices.
BERT (Bidirectional Encoder Representations from Transformers) and GPT (Generative Pre-trained Transformer) are two landmark architectures in the development of large language models (LLMs). BERT, introduced by Google, uses a bidirectional training approach, allowing it to capture context from both directions of a sentence simultaneously. This makes it particularly effective for tasks like sentence classification, question answering, and named entity recognition, where understanding relationships between words is crucial [3]. On the other hand, GPT, developed by OpenAI, takes an autoregressive approach, generating text word by word in a unidirectional manner. Starting with GPT-1 and evolving through GPT-4, GPT has proven its strength in creative and coherent text generation, conversational AI, and text summarization, making it a highly flexible model for both content generation and interactive applications [8, 5].
More recent models like Meta’s LLaMA (Large Language Model Meta AI) and Gemma are pushing the boundaries of LLM development. LLaMA is designed to be highly efficient in terms of computational resources, offering a smaller model that still performs well on a wide range of natural language processing tasks. Its streamlined design makes it accessible for research and industrial applications where resources might be limited [6]. Gemma, an emerging model family, is expected to further refine the efficiency and adaptability of LLMs, with an emphasis on real-time performance and scalability. Gemma is particularly well-suited for industry-specific applications, aiming to bridge the gap between general-purpose LLMs and more specialized use cases [7]. Together, these model families reflect the rapid progression and specialization in the field of large language models, each offering distinct strengths based on the needs of users and developers.
@@
This study introduces an innovative attention layer designed to enhance the efficiency of inference stages in Transformer-based models with an adjustable approach. The suggested solution involves a sorting and elimination process for hidden sequence elements (word-vectors), resulting in a reduction of effective word-vectors in each layer and a decrease in the number of FLOPs. The primary contributions of this research can be summarized as follows:
@@
  • An accurate analytical inference speedup estimation is presented for the proposed approach that helps a system designer determine the hyper-parameter selection at the fine-tuning phase.
@@
  • The proposed approach can be applied to a variety of Transformer-based architectures, while preserving the global context of the input.
@@ II-B Head/Channel Pruning
A parameter reduction framework was proposed by identifying the most critical heads in each encoder layer using layer-wise relevance propagation and pruning the redundant heads [35]. It showed that the earlier encoder layer’s heads are much more critical than the last. Moreover, in [36], it is shown that around 17% of attention heads can be removed at the test time without significantly impacting inference performance. In [37], a rapid post-training pruning framework tailored for Transformers is introduced. This framework involves three key fine-tuning stages: mask search, mask rearrangement, and mask tuning. Notably, the method obviates the necessity for any retraining, presenting an efficient approach to achieve model compression in the Transformer architecture.
@@ II-E Non-Pruning Methods
ALBERT is a lighter version of BERT model [48]. It uses cross-layer weight sharing and a modified embedding layer structure, making the embedding size independent of the encoder’s hidden size. Moreover, the new Sentence Order Prediction (SOP) task is utilized in the pre-training phase instead of the Next Sentence Prediction task. ALBERT reports state-of-the-art results with its parameter reduction and modified pre-training policy. ALBERT is nine times smaller and 1.2 times faster than BERT-base, with less than 2.5% on average accuracy drop. ALBERT is an example of the parameter-sharing technique.
DistilBERT [49], TinyBERT [50] and MobileBERT [51] use the knowledge distillation technique during the pre-training and offer up to 5.5 times faster inference on downstream tasks than BERT-base. In [52] a Self-Distillation (SD) mechanism obtains high-accuracy models directly without going through an assistive model. The main concerns with these techniques are the pre-training effort of the model from scratch, extra training steps, and significant accuracy degradation.
In [53], the authors propose sufficient conditions under which they prove that a sparse attention model can universally approximate any sequence-to-sequence function. They show that sparse Transformers with only $\mathcal{O}_{n}$ connections per attention layer can approximate the same function class as the dense model with $n^{2}$ connections. Consequently, in recent years, many works [54, 55, 56, 57, 58, 59, 60, 61, 62] have proposed methods that aim to reduce the complexity of the attention mechanism.
@@
The previous works are trying to reduce the model size and inference latency. These works show that significant part computations in the models appear unnecessary, and models with less computational complexity can be obtained from the original ones. The main drawbacks of these methods are pre-training from scratch, multi-phase fine-tuning, and significant inference accuracy drop. In these methods, the inference speedup is set in the fine-tuning phase, and the achieved speedup is fixed in the inference phase. The proposed solution tries to mitigate the previous drawbacks. It presents the powerful offline-tuning mechanism that controls the inference speedup of the model after the fine-tuning phase.
The impact of the speedup methods on the model stability and the input’s global context understanding is an important concern that must be studied in solutions evaluations. Backpack learns multiple non-contextual sense vectors for each word in a vocabulary and represents a word in a sequence as a context-dependent, non-negative linear combination of sense vectors in this sequence [66]. The Backpack Model shows that each word in the input sequence can interpreted as a sense vector that encodes different aspects of a word which is very sensitive. A good practice for assessing the performance of a method is evaluating it in a more complex task, like text generation tasks that inspect the model proficiency. Therefore the suggested method is applied to the several autoregressive models and evaluated on several text generation tasks.
Caching methods focus on optimizing the efficiency and speed of inference in autoregressive large language models (LLMs). KV-Runahead enhances prompt phase parallelization by using multiple processes to manage KV-cache population, thereby reducing the Time-to-First-Token (TTFT) [67]. CacheGen speeds up context loading with a custom tensor encoder that compresses KV-cache into compact bitstream representations, minimizing decoding overhead [68]. CacheBlend introduces a method for reusing pre-computed KV-caches and selectively updating a small subset of tokens, optimizing cache reuse [69]. Prompt Cache accelerates inference by precomputing and storing attention states for frequently occurring text segments, enabling efficient reuse across different prompts [70]. Lastly, CachedAttention employs a hierarchical KV-caching system using cost-effective memory and storage solutions to manage KV-caches for all requests [71]. Together, these methods contribute to faster and more efficient LLM inference by improving cache management and state reuse.
@@ III Proposed Method
This section presents details of the proposed method and the analysis and reasoning behind the approach. Section III-A introduces the preliminary background of the Transformer architecture, while Section III-B analyzes the latency and TTFT across various Transformer-based models. The proposed encoder layer is based on an insight into the amount of the contribution of word-vectors at different layers, which is considered in Section III-C. Section III-D explores the proposed architecture in detail. Additionally, Section III-E extends the proposed method to autoregressive architectures and applies it to several newly developed models.
@@ III-A Background
Despite the current progress and advancements in mainstream LLM architectures such as GPT, Gemma, Mistral [72], and Llama, these models still follow the foundational Transformer architecture [1]. The primary differences among them lie in the number of parameters, layers, heads, hidden states, and specific modifications to normalization layers, residual paths, and attention implementations. For clarity and a deeper understanding of the suggested method, we first focus on the BERT [48] and GPT-2 [73] models before extending the approach to other large language models (LLMs).
Fig. 1 shows BERT-base model architecture and output classifier. The input sentence goes through the tokenizer layer and embedding layer. Then the context is fed into 12 cascaded encoder layers to be processed. The pooler and classifier layers are added in the last layer to perform proper output for the desired task. In the figure, $T$ and $H$ are the numbers of input word-vectors and the hidden state vector size, respectively. Each encoder consists of a Bidirectional Multi-Head Self-Attention layer and a Feed-Forward layer with several tensor multiplications. The residual branches and normalization layers increase the training stability.
Figure 1: BERT-base [3] architecture with output layers
III-B Transformer Architecture Latency and TTFT Analysis
Table I presents an analysis of the latency and TTFT share across the embedding, attention, and feed-forward layers in several Transformer-based models. As shown in the table, regardless of model size, the latency of the embedding layer is minimal, with the majority of latency attributed to the attention and feed-forward layers. Attention share ranges from around 20% to 47% of the total. In neural networks, latency and computational effort are closely related to the number of FLOPs. Consequently, reducing the number of FLOPs in the encoder layers, which account for most of the latency, significantly decreases both total latency and TTFT.
TABLE I: Latency and Time-to-First-Token (TTFT) Analysis for BERT [3], GPT-2 [73], T5 [74], Gemma2 [75],
Mistral [72], and Llama3 [76] Models
Model Architecture Type Parameters Layers Attention Heads Latency/TTFT Sharea (%)
Embedding Attention Feed-Forward
BERT-base Encoder 110M 12 12 0.58 34.94 64.48
GPT-2 Decoder 124M 12 12 0.06 34.94 65.00
T5-base Encoder-Decoder 220M 12 12 0.35 47.00 52.65
T5-large Encoder-Decoder 770M 24 16 0.06 40.46 59.47
GPT-2-large Decoder 774M 36 20 0.01 32.25 67.74
Gemma2 Decoder 2B 26 8 0.01 20.36 79.63
Mistral Decoder 7B 32 32 0.01 20.34 79.65
Llama3 Decoder 8B 32 32 0.01 19.55 80.44
a Latency share for BERT model and Time-to-First-Token (TTFT) for other autoregressive models
@@ III-C Attention Context Contribution (ACC) metric
Fig. 2 illustrates the calculation of the Score Vector based on the attention probability matrix. Fig. 2 (left) shows the attention probability matrix of BERT-base model. The rows of this matrix are normalized to one. By averaging the attention probability matrix over the heads, the aggregated probability matrix is resulted, Fig. 2 (center). The SV results from the sum of this matrix along the unnormalized dimension as shown in Fig. 2 (right). The SV elements with higher score values correspond to the word-vectors that contribute more to the layer output.
@@
Fig. 3 illustrates BERT-base SV results on the IMDB sentiment analysis task[77] for an input sample. As shown in earlier encoder layers, the SV values approximately have uniform distributions, and in the last encoder layers, the distribution of SV values changes to approximately $\delta$ distribution. The distribution of SV values in the last layers indicates that the concentration of scores is high on a few word-vectors and others have less contribution to the self-attention layer output.
Figure 3: BERT-base[3] Score Vector (SV) results on the IMDB sentiment analysis task[77]
The SV elements are positive and with an average of one. So, the median of SV is bounded. Suppose the median of the SV is lower than its average. In that case, more than half of the elements are lower than the average value, and only a few are higher than the average value. It suggests that most of the word-vectors do not bear enough contribution to the layer output, and a few word-vectors influence the final result of the layer. The median of the Score Vector is interpreted as the proposed Attention Context Contribution metric.
The ACC metric calculation is explained in Algorithm 1. The algorithm inputs are BERT encoder word-vectors and the model parameters. The algorithm output is the ACC metric vector that suggests the contribution of the word-vectors in each encoder self-attention layer. For encoder layer $l$, the first step is to prepare the pre-computed self-attention probability matrix; $M_{Att}^{l}$. Then the average of this matrix over its heads is computed; $M_{H}^{l}$. In the next step, the resulting matrix’s summation over its unnormalized dimension is computed; $M_{S}^{l}$. This matrix is the Score Vector. Attention Context Contribution of the layer ($E_{ACC}^{l}$) is the median of the Score Vector.
@@
Fig. 4 shows the ACC metric for BERT-base encoder layers on IMDB [77] dataset and several General Language Understanding Evaluation (GLUE) [78] benchmark tasks. The fitted curve to the ACC metric results shows that the ACC metric is reduced at later layers gradually. It indicates that the fraction of the word-vectors that contribute more in the last layers is less in those layers; hence more word-vectors can be eliminated. By eliminating less important word-vectors in each layer, the processed word-vectors are decreased gradually, and the inference time of the model is decreased.
@@
Figure 4: BERT-base [3] Encoder Attention Context Contribution (ACC) metric results (blue bars) and second degree fitted curves (red lines) on SST-2 [79], MRPC [80], QQP [81], QNLI [82], RTE [83], MNLI [84], STS-B [85] and IMDB [77] tasks.